Nerdyhuman:

What is the period of the wave motion for a wave with a frequency of 0.37 kHz?

Nerdyhuman:

@vocaloid

Vocaloid:

period = 1/frequency if you want SI units, convert KHz to Hz first. your solution will be in seconds.

Nerdyhuman:

A sound wave has a frequency of 2004 Hz. What is the distance between crests or compressions of the wave? (Take the speed of sound to be 344 m/s.) Thank you! I just need formulas relly

Vocaloid:

wave speed = frequency * wavelength ("distance between crests or compressions" is another way to say wavelength)

Nerdyhuman:

Compute the wavelength of an X-ray with a frequency of 8.0 1018 Hz. Got it!

Vocaloid:

same formula, but for electromagnetic radiation like x-rays you'll use c = speed of light

Nerdyhuman:

Will it be in nm or will I need to convert? I have 3.75 but it says it is wrong

Vocaloid:

notice how you're in SI units (m/s) which means your answer will be in m if it asks for nm you need to convert

Nerdyhuman:

so the example basically moved the answer into a decimal but I put .375 and that is wrong

Vocaloid:

c = wavelength * frequency 299,792,458 = (x) * (8.0 * 10^18) x = 3.74741×10^-11 m = 0.0375 nm

Nerdyhuman:

oh I forgot a decimal place okay cool

Nerdyhuman:

Picnickers see a lightning flash and hear the resulting thunder 7.60 s later. If the storm is traveling at a rate of 13.0 km/h, how long, in minutes, do the picnickers have before the storm arrives at their location? I have like 3 more questions but thank you so much for your help

Vocaloid:

not sure about this one. has something to do with how fast sound travels, and d = rt.

Vocaloid:

did a quick search of this problem from the moment the lightening strikes, the sound has 7.60 seconds to travel. sound travels 340 m/s, which means the sound travelled 340 * 7.60 = 2584 m and thus thus the storm is about 2584 meters away. the storm travels 13.0km/hr, so d = rt 2584m = (13000m/hr) * t t = 0.19877 hours = 11.926 minutes

Nerdyhuman:

During a thunderstorm, 5.6 s elapses between observing a lightning flash and hearing the resulting thunder. Approximately how far away in kilometers and miles was the lightning flash? (Assume the speed of sound is 344 m/s.) Is this similar to the previous problem?

Vocaloid:

I think so, I think you just need to multiply speed of sound by 5.6 s

Nerdyhuman:

that's not it it said it was wrong

Vocaloid:

it says km, did you remember to divide by 1000

Vocaloid:

it also asks for miles so you also need to do a km --> miles conversion

Nerdyhuman:

so I got the 1.926 correct for the kilometers and tried to convert and got it wrong for miles I got 1926 for the miles

Vocaloid:

keep in mind it says miles, not meters 1 km = 0.6214 miles, so 1.926km = 1.1968 miles

Nerdyhuman:

ooooooh right okay

Nerdyhuman:

(a) Approximately how long would it take a telephone signal to travel 2800 mi from coast to coast across the United States? (Telephone signals travel at about the speed of light.) one last question this has a part a and b

Vocaloid:

simple d = rt problem 2800 miles = 4.506 * 10^6 meters speed = the speed of light = 299,792,458 m/s plugging into d = rt ---> 4.506 * 10^6 meters = 299,792,458 m/s * (time) solve for time. that'll give you time in seconds. convert to whatever unit they ask for.

Nerdyhuman:

thank you! all done